OpenStudy (anonymous):

The stopping distance d of a car after the brakes are applied vaires directly as the square of the speed r. If a car traveling 70 mph can stop in 270 ft, how many feet will it take the same car to stop when it is traveling 80 mph?

OpenStudy (anonymous):

notice that your problem says that the distance (d) varies directly as the square of the speed r. This means that basically we have f(d)=r^2 Which means that our problem becomes a simple case of cross multiplication since we have: 70^2/270=80^2/x You can work that out, it will give you the answer

OpenStudy (anonymous):

First the relationship is such that the stopping distance varies directly with square of speed ie. D(distance) = K (proportionality constant) * r^2 (speed) The units in the given relationships are different to each other….since the question is asking us to calculate how many feet…..lets first convert the miles per hour into ft per sec 70 miles/hr=1026.36 ft/s (verify yourself) 80 miles/ hr=1172.98ft/s Now lets use the given condition to find the constant ‘K’ D=K*r^2 270 =K*(1026.36)^2 K=.000256 Plug the value of K to find the stopping distance when the speed is 80 miles/hr ie. 1172.98ft/s D=.000256*( 1172.98)^2 D=352.22 ft

OpenStudy (anonymous):

it not that complicated